INFRARED SPECTRA AND INTENSITIES OF WATER COMPLEXES WITH NITROGEN, OXYGEN AND ARGON IN HELIUM DROPLETS

\underline{SUSUMU KUMA}, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; MIKHAIL N. SLIPCHENKO^{a}, KIRILL E. KUYANOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90080, USA; TAKAMASA MOMOSE, Department of Chemistry, University of British Columbia, Vancouver, BC V6T 1Z1, Canada; and AN-DREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90080, USA.

The infrared spectra of N_{2}-H_{2}O, O_{2}-H_{2}O and Ar-H_{2}O complexes in superfluid He droplets were measured in the range of the stretching vibrational bands of water molecules. The infrared intensity of the anti-symmetric stretching bands in these complexes were found to be very similar to that in single H_{2}O molecules. The spectra show that H_{2}O in O_{2}-H_{2}O and Ar-H_{2}O rotates nearly freely, whereas no indication of H_{2}O internal rotation was observed in N_{2}-H_{2}O spectra. The conformation of the N_{2}-H_{2}O complexes was estimated from the rotational constants.

apresent address:1605 Gilman Hall, Iowa State University, Ames, IA 50011-3111, USA